# Enhanced Fish Freshness Classification with Incremental Handcrafted Feature Fusion

### v1

@@ Abstract
Accurate assessment of fish freshness remains a major challenge in the food industry, with direct consequences for product quality, market value, and consumer health. Conventional sensory evaluation is inherently subjective, inconsistent, and difficult to standardize across contexts, often limited by subtle, species-dependent spoilage cues. To address these limitations, we propose a handcrafted feature–based approach that systematically extracts and incrementally fuses complementary descriptors, including color statistics, histograms across multiple color spaces, and texture features such as Local Binary Patterns (LBP) and Gray-Level Co-occurrence Matrices (GLCM), from fish eye images. Our method captures global chromatic variations from full images and localized degradations from ROI segments, fusing each independently to evaluate their effectiveness in assessing freshness. Experiments on the Freshness of the Fish Eyes (FFE) dataset demonstrate the approach’s effectiveness: in a standard train–test setting, a LightGBM classifier achieved 77.56% accuracy, a 14.35% improvement over the previous deep learning baseline of 63.21%. With augmented data, an Artificial Neural Network (ANN) reached 97.16% accuracy, surpassing the prior best of 77.3% by 19.86%. These results demonstrate that carefully engineered, handcrafted features, when strategically processed, yield a robust, interpretable, and reliable solution for automated fish freshness assessment, providing valuable insights for practical applications in food quality monitoring.
@@
organization=Department of Artificial Intelligence, FPT University, addressline=Da Nang, postcode=550000, country=Viet Nam
@@ 3.1.1 Color features
In addition, percentile values (such as the $5^{\text{th}}$, $25^{\text{th}}$, $50^{\text{th}}$, $75^{\text{th}}$, and $95^{\text{th}}$) are computed for each component of the color spaces to capture the distribution of color intensities, providing more insight into the image’s color characteristics. Fifteen features are extracted per color space (five percentiles × three components).
@@ 3.1.2 Texture features
Gray Level Co-occurrence Matrix (GLCM): GLCM is a statistical method for analyzing texture based on the spatial relationship between pixel pairs. In this study, the GLCM is computed only for the b* channel in the Lab color space at a fixed pixel distance (d = 3) and four orientations (0°, 45°, 90°, and 135°). From each GLCM, a set of Haralick features, including contrast, homogeneity, energy, and correlation, is extracted [29]. These features are then aggregated across the four orientations using statistical measures such as the mean and range, resulting in a compact and rotation-invariant texture descriptor specific to the b* channel.
@@ 3.2 Eye region localization and segmentation
This section outlines the procedure for segmenting the eye region from fish images, forming the basis for localized feature extraction. To assess the effect of isolating the region of interest (ROI), we developed an automatic segmentation method based on radial intensity analysis. This algorithm efficiently detects eye boundaries without relying on deep learning models. The procedure consists of four sequential stages:
Step 1 (Image preprocessing): The original color image is converted to grayscale to simplify intensity analysis. A Gaussian filter with a $7\times 7$ kernel is then applied to suppress noise and irrelevant fine details, thereby enhancing salient structures and stabilizing boundary detection. The image center, defined as the midpoint of its width and height, is assumed to coincide with the eye center.
Step 2 (Radial intensity scanning for boundary detection): From the defined center, the algorithm emits uniformly spaced radial scan lines over $360^{\circ}$. Along each line, pixel intensities are sampled from the center outward, producing a one-dimensional profile of intensity variation. The first-order gradient of this profile is then computed to capture brightness changes. The eye boundary, marked by a transition from the dark pupil/iris to the brighter surrounding tissue, is detected as the location with the steepest negative gradient. This point on each radial line is recorded as a candidate boundary point.
Step 3 (Radius estimation and outlier removal): The set of candidate radii obtained from all radial scans is aggregated, representing distances from the center to detected boundary points. Outliers caused by specular reflections on the moist, curved eye surface are removed by using the median, which provides a robust estimate against extreme values. The final eye radius is derived from this median and slightly adjusted to ensure complete coverage of the eye region.
Step 4 (Mask generation and segmentation): Based on the estimated eye center and final radius, a binary mask is generated. The mask has the same dimensions as the original image, with a filled white circle denoting the segmented eye region against a black background. Applying this mask to the original color image via a logical AND operation preserves only the eye region while setting all surrounding pixels to black. This segmented output is then used for subsequent feature extraction. An example of the segmentation result is illustrated in Figure 2.
Figure 2: Comparison of eye images before and after segmentation
@@ 4.1 Data description
The dataset employed in this study is the publicly available the Freshness of the Fish Eyes (FFE) dataset, originally introduced and described by Prasetyo et al. [9], and accessible through Mendeley data [38]. The FFE dataset consists of sRGB images of fish eyes captured to evaluate different levels of freshness. It includes images from eight fish species, with each image labeled into one of three freshness categories: ‘Highly fresh’ (corresponding to a 1-2 day shelf life), ‘Fresh’ (3–4 days), and ‘Not fresh’ (5-6 days). In total, the dataset comprises 4,390 images, with representative examples shown in Figure 3.
In this study, we adopted a generalized freshness classification strategy. Specifically, instead of distinguishing among the eight fish species, we pooled all images based on their assigned freshness levels. This simplification yields a three-class problem that directly captures the visual indicators of freshness observable across species. Accordingly, the dataset used in our experiments is distributed into three categories: 1764 images labeled as highly fresh, 1320 as fresh, and 1306 as not fresh. The dataset was stratified and initially divided into 80% for training and 20 for testing. The training subset was then further stratified and split into 80% for training and 20% for validation, as illustrated in Figure 4.
@@ 4.3 Experiment setting
ANN activation=tanh, solver=Adam, learning_rate=adaptive, early_stopping=False, max_iter=500, alpha=0.001, hidden_layer_sizes=(128)
@@ 5.2 Results on segmented images
To examine whether focusing only on the eye, the primary region of interest, could affect classification performance, we evaluated the impact of isolating the fish eye using segmented images. By removing the surrounding background and tissues, we aimed to examine whether the models could better focus on the visual cues of freshness in the eye itself and to assess the importance of the contextual information from the surrounding regions. The results, summarized in Table 9, reveal that the surrounding tissues and background provide complementary cues that are critical for accurate freshness classification.
@@ 5.3 Computation time analysis
These findings highlight the trade-off between computational cost and classification performance. As shown in Table 4, CIELAB CS achieved the highest standalone accuracy (70.96%) despite its heavy computation. Its inclusion in the final best-performing model (FS17) confirms its critical role in maximizing accuracy. Overall, the computation time analysis not only demonstrates this accuracy–speed trade-off but also provides a practical framework for balancing performance and efficiency in real-world deployment.
@@ 5.4.1 Re-implementation of prior methodology
To ensure a fair comparison with previous studies, we first re-implemented the experimental settings described in Yildiz et al. [10], including their data augmentation strategy and 10-folds cross-validation. All hyperparameters of the classifiers were kept identical to the original study, allowing a direct assessment of the impact of feature representation without introducing confounding factors. Moreover, for each classifier, we used the reported best-performing results from Yildiz et al. to establish a conservative and fair baseline. This ensures that any observed improvements in performance can be attributed solely to the proposed handcrafted feature extraction rather than differences in data handling, parameter tuning, or selection bias.
The results presented in Table 11 clearly demonstrate the substantial advantage of the proposed handcrafted feature extraction approach over deep feature representations (VGG19) across multiple ML classifiers. Notably, ANN achieves an accuracy of 97.16%, outperforming the corresponding VGG19-based model by nearly 20%. This significant improvement highlights the effectiveness of the carefully designed feature set, comprising statistical color moments, color histograms, LBP, and GLCM, in capturing the visual cues most relevant to fish eye freshness classification.
Other classifiers further reinforce this trend. For instance, RF attains 92.06% accuracy, surpassing the VGG19-based counterpart by nearly 24 points, and KNN also benefits from a +5% gain. Even linear models such as LR and SVM perform comparably or slightly better than VGG19 features, demonstrating that the handcrafted features are robust across different learning paradigms. Moreover, ensemble methods like ET, LGBM, and CB reach accuracy levels above 96%, emphasizing that the proposed features provide discriminative information that can be fully leveraged by tree-based models without requiring deep feature extraction.
These results provide compelling evidence that domain-specific, handcrafted features can not only compete with but often surpass pretrained deep features, even when comparing against the best results reported in previous studies. This underscores the value of careful feature engineering in achieving high-performance and reliable classification in specialized visual inspection tasks, offering a practical and effective alternative to deep learning approaches for similar datasets.
Table 11: Comparison of ML model accuracy (%) between the proposed handcrafted features and deep features reported by Yildiz et al. [10].
@@
KNN 72.76 67.8 4.96
LR 63.62 65.9 -2.28
SVM 73.25 74.3 -1.05
ANN 97.16 77.3 19.86
RF 92.06 68.3 23.76
ET 96.91 - -
LGBM 97.53 - -
CB 97.23 - -
5.4.2 Comparative analysis on the non-augmented dataset
In addition to re-implementing the experimental setup with data augmentation from Yildiz et al. [10], we also evaluated our approach using the standard train-test split proposed by Prasetyo et al. [9], who reported a baseline accuracy of 63.21% with their custom deep learning architecture, MB-BE (a modified MobileNetV1), on the original, non-augmented dataset.
@@ 6 Conclusion
This study proposed and systematically evaluated a handcrafted feature engineering approach for fish freshness classification from eye images. We extracted a diverse set of descriptors, including color statistics, variance ratios, percentiles, histograms, and texture features (GLCM and LBP), and integrated them through an incremental fusion strategy to form a rich numerical representation of freshness cues. The extracted features were subsequently used to train and evaluate eight ML models—LR, KNN, SVM, ANN, RF, ET, LGBM, and CB—to comprehensively assess classification performance. Experiments on the FFE dataset demonstrated that, in the standard train-test setting, our LGBM model achieved 77.56% accuracy, an improvement of over 14% compared to the 63.21% reported by Prasetyo et al. [9]. Under an augmented data setting, our ANN model further reached 97.16%, surpassing the 77.3% reported by Yildiz et al. [10], thereby establishing a new state-of-the-art benchmark. Importantly, comparative analyses between full images and ROI-based segmentation revealed that contextual information from surrounding tissues contributes significantly to classification performance, suggesting that freshness cues extend beyond the eye region alone.
The main limitation of this approach lies in its reliance on handcrafted features, which require significant domain expertise and may not generalize well to other fish species or imaging conditions. Moreover, these features might miss abstract patterns that deep learning models could capture automatically. Future work will address this by applying advanced deep learning models and hybrid approaches that integrate handcrafted features with deep features, combining interpretability with the ability to learn additional representations for more robust and accurate fish freshness assessment.

### v2

@@ Abstract
Accurate assessment of fish freshness remains a major challenge in the food industry, with direct consequences for product quality, market value, and consumer health. Conventional sensory evaluation is inherently subjective, inconsistent, and difficult to standardize across contexts, often limited by subtle, species-dependent spoilage cues. To address these limitations, we propose a handcrafted feature–based approach that systematically extracts and incrementally fuses complementary descriptors, including color statistics, histograms across multiple color spaces, and texture features such as Local Binary Patterns (LBP) and Gray-Level Co-occurrence Matrices (GLCM), from fish eye images. Our method captures global chromatic variations from full images and localized degradations from ROI segments, fusing each independently to evaluate their effectiveness in assessing freshness. Experiments on the Freshness of the Fish Eyes (FFE) dataset demonstrate the approach’s effectiveness: in a standard train-test setting, a LightGBM classifier achieved 77.56% accuracy, a 14.35% improvement over the previous deep learning baseline of 63.21%. With augmented data, an Artificial Neural Network (ANN) reached 97.49% accuracy, surpassing the prior best of 77.3% by 20.19%. These results demonstrate that carefully engineered, handcrafted features, when strategically processed, yield a robust, interpretable, and reliable solution for automated fish freshness assessment, providing valuable insights for practical applications in food quality monitoring.
@@
organization=Faculty of Artificial Intelligence,
FPT University, addressline=Da Nang, postcode=550000, country=Viet Nam
@@ 3.1.1 Color features
In addition, percentile values (such as the $5^{\text{th}}$, $25^{\text{th}}$, $50^{\text{th}}$, $75^{\text{th}}$, and $95^{\text{th}}$) are computed for each component of the color spaces to capture the distribution of color intensities, providing more insight into the image’s color characteristics. Fifteen features are extracted per color space (five percentiles $\times$ three components).
@@ 3.1.2 Texture features
Gray Level Co-occurrence Matrix (GLCM): GLCM is a statistical method for analyzing texture based on the spatial relationship between pixel pairs. In this study, the GLCM is computed only for the b* channel in the Lab color space at a fixed pixel distance (d = 3) and four orientations ($0^{\circ}$, $45^{\circ}$, $90^{\circ}$, and $135^{\circ}$). From each GLCM, a set of Haralick features, including contrast, homogeneity, energy, and correlation, is extracted [29]. These features are then aggregated across the four orientations using statistical measures such as the mean and range, resulting in a compact and rotation-invariant texture descriptor specific to the b* channel.
@@ 3.2 Eye region localization and segmentation
This section outlines the procedure for segmenting the eye region from fish images, forming the basis for localized feature extraction. To isolate the region of interest (ROI) and eliminate background noise, we developed an automatic segmentation algorithm based on radial intensity analysis, which detects eye boundaries without relying on deep learning models. Let $I(x,y)$ denote the original RGB input image of size $W\times H$. The overall segmentation process consists of four sequential stages:
Step 1 (Image preprocessing): First, the color image $I$ is converted to a grayscale intensity image $I_{\mathrm{gray}}$. To suppress high-frequency noise and stabilize the subsequent gradient calculation, a Gaussian filter $G$ with a kernel size of $7\times 7$ is convolved with the image:
$I_{\mathrm{smooth}}(x,y)=I_{\mathrm{gray}}(x,y)\ast G_{7\times 7}(x,y).$ (1)
The scanning center is defined as the geometric center of the image:
$(c_{x},c_{y})=\left(\frac{W}{2},\frac{H}{2}\right).$ (2)
Step 2 (Radial intensity scanning for boundary detection): Radial scan lines are then emitted from the image center at angles $\theta\in[0,360^{\circ})$ with a step of $2^{\circ}$. For each direction $\theta$, the intensity profile $P_{\theta}(r)$ is sampled at a distance $r$ from the center:
$x_{r}$ $=c_{x}+r\cos(\theta),$ (3)
$y_{r}$ $=c_{y}+r\sin(\theta),$
where $P_{\theta}(r)=I_{\mathrm{smooth}}(x_{r},y_{r})$. To detect the eye boundary, we compute the discrete gradient of the intensity profile:
$\nabla P_{\theta}(r)=P_{\theta}(r+1)-P_{\theta}(r).$ (4)
Since the eye region is typically brighter than the surrounding skin/scales, the boundary is characterized by a significant drop in intensity (negative gradient). The candidate radius $r_{\theta}^{*}$ is identified at the point of the steepest negative gradient:
$r_{\theta}^{*}=\arg\min_{r}\,\nabla P_{\theta}(r),\quad\text{subject to }\nabla P_{\theta}(r)<\tau,$ (5)
where $\tau=-5$ is a threshold to filter out weak edges.
Step 3 (Radius estimation and outlier removal): Scanning across all angles yields a set of candidate radii $\mathcal{R}=\{r_{\theta}^{*}\}$. To eliminate outliers caused by specular reflections, we compute the median of the set. The final radius $R_{\mathrm{final}}$ is then expanded by a factor $\lambda=1.25$ to ensure the entire eye region is covered:
$R_{\mathrm{final}}=\left\lfloor\lambda\cdot\mathrm{median}(\mathcal{R})\right\rfloor,$ (6)
Step 4 (Mask generation and segmentation): Finally, a binary mask $M(x,y)$ is generated using the estimated center and expanded radius:
$M(x,y)=\begin{cases}1,&\text{if }\sqrt{(x-c_{x})^{2}+(y-c_{y})^{2}}\leq R_{\mathrm{final}},\\ 0,&\text{otherwise}.\end{cases}$ (7)
The segmented eye image $I_{ROI}$ is obtained by applying the mask to the original image via a bitwise AND operation:
$I_{\mathrm{ROI}}(x,y)=I(x,y)\cdot M(x,y).$ (8)
An example of the segmentation result is illustrated in Figure 2.
Figure 2: Comparison of eye images before and after segmentation.
@@ 4.1 Data description
The dataset employed in this study is the publicly available the Freshness of the Fish Eyes (FFE) dataset, originally introduced and described by Prasetyo et al. [9], and accessible through Mendeley data [38]. The FFE dataset consists of sRGB images of fish eyes captured to evaluate different levels of freshness. It includes images from eight fish species, with each image labeled into one of three freshness categories: ‘Highly fresh’ (corresponding to a 1–2 day shelf life), ‘Fresh’ (3–4 days), and ‘Not fresh’ (5–6 days). In total, the dataset comprises 4,390 images, with representative examples shown in Figure 3.
In this study, we adopted a generalized freshness classification strategy. Specifically, instead of distinguishing among the eight fish species, we pooled all images based on their assigned freshness levels. This simplification yields a three-class problem that directly captures the visual indicators of freshness observable across species. Accordingly, the dataset used in our experiments is distributed into three categories: 1764 images labeled as highly fresh, 1320 as fresh, and 1306 as not fresh. The dataset was stratified and initially divided into 80% for training and 20% for testing. The training subset was then further stratified and split into 80% for training and 20% for validation, as illustrated in Figure 4.
@@
4.2 Evaluation protocols
To ensure reliable generalization assessment and fair comparison with prior studies, two complementary evaluation protocols were adopted in this work.
Protocol A (PA – hold-out evaluation)
Protocol A is used as the primary evaluation setting to assess the effectiveness of the proposed handcrafted feature extraction and incremental fusion strategy under a standard hold-out evaluation scheme. This protocol follows the stratified train–test split strategy adopted in earlier foundational studies on the FFE dataset [9]. Specifically, the dataset is split into training and test subsets using a stratified hold-out scheme to preserve class distributions. All models are trained and evaluated solely on the original images, ensuring that each test sample is completely unseen during training. Results obtained under Protocol A reflect the generalization performance of the proposed method and form the basis for all main analyses and conclusions in this paper.
Protocol B (PB – replication setting)
Protocol B follows the experimental configuration established in previous studies [10] to facilitate a direct comparison. In this setting, data augmentation (consisting of image flipping) is performed prior to dataset splitting, and the evaluation is conducted using 10-fold cross-validation. This protocol is specifically adopted to replicate reported results and to evaluate the proposed handcrafted feature fusion method under identical experimental conditions. Given that augmentation precedes splitting, results obtained under Protocol B should be interpreted as upper-bound performance within the original experimental framework rather than a measure of real-world generalization.
@@ 4.4 Experiment setting
[…]
@@ 5.1 Results on full images
All results reported in this subsection are obtained under Protocol A (PA), following a clean hold-out evaluation without data augmentation.
@@ 5.2 Results on segmented images
To investigate whether focusing exclusively on the eye-the primary region of interest-affects classification performance, we evaluated the proposed approach using segmented eye images under Protocol A, enabling a direct comparison with the corresponding results obtained from full images. By removing the surrounding background and tissues, we aimed to examine whether the models could better focus on the visual cues of freshness in the eye itself and to assess the importance of the contextual information from the surrounding regions. The results, summarized in Table 9, reveal that the surrounding tissues and background provide complementary cues that are critical for accurate freshness classification.
@@ 5.3 Computation time analysis
These findings highlight the trade-off between computational cost and classification performance. As shown in Table 4, CIELAB CS achieved the highest standalone accuracy (70.96%) despite its heavy computation. Its inclusion in the final best-performing model (FS17) confirms its critical role in maximizing accuracy. Overall, the computation time analysis not only demonstrates this accuracy-speed trade-off but also provides a practical framework for balancing performance and efficiency in real-world deployment.
@@ 5.4 Comparison with previous studies on the same dataset
To ensure a fair and faithful comparison with prior work, this section reports results obtained under two different evaluation settings: an augmented data setting combined with cross-validation and a hold-out evaluation on the original, non-augmented dataset, following the experimental protocol adopted in previous studies.
@@ 5.4.1 Re-implementation of prior methodology
In this subsection, experiments are conducted under the augmented data and cross-validation setting adopted in previous studies (Protocol B). To ensure a fair and faithful comparison with prior work, we re-implemented the experimental protocol described by Yildiz et al. [10], including their strategy of applying mirroring-based data augmentation to the entire dataset before performing 10-fold cross-validation. The hyperparameters of the classifiers inherited from the original study (KNN, LR, SVM, ANN, and RF) were kept identical to those reported in [10]. For the additional models introduced in this work (ET, LGBM, and CB), appropriate hyperparameter settings were selected. The complete configurations of all classifiers are summarized in Table 3.
The results presented in Table 11 clearly demonstrate the substantial advantage of the proposed handcrafted feature extraction approach over deep feature representations (VGG19, selected as the strongest baseline in [10]) across multiple ML classifiers. Notably, ANN achieves an accuracy of 97.49%, outperforming the corresponding VGG19-based model by over 20%. This significant improvement highlights the effectiveness of the carefully designed feature set, comprising statistical color moments, color histograms, LBP, and GLCM, in capturing the visual cues most relevant to fish eye freshness classification.
Other classifiers further reinforce this trend. For instance, RF attains 91.91% accuracy, surpassing the VGG19-based counterpart by nearly 24 points, and KNN also benefits from a gain of about 5%. Even linear models such as LR and SVM perform comparably or slightly better than VGG19 features, demonstrating that the handcrafted features are robust across different learning paradigms. Moreover, ensemble methods like ET, LGBM, and CB reach accuracy levels above 97%, emphasizing that the proposed features provide discriminative information that can be fully leveraged by tree-based models without requiring deep feature extraction.
Table 11: Comparison of ML model accuracy (mean $\pm$ standard deviation %) between the proposed handcrafted features and deep features reported by Yildiz et al. [10].
@@
KNN $72.88\pm 1.48$ 67.8 $5.08\pm 1.48$
LR $63.14\pm 1.78$ 65.9 $2.76\pm 1.78$
SVM $73.23\pm 1.88$ 74.3 $1.07\pm 1.88$
ANN $\mathbf{97.49\pm 0.63}$ 77.3 $20.19\pm 0.63$
RF $91.91\pm 1.16$ 68.3 $23.61\pm 1.16$
ET $\mathbf{97.16\pm 0.88}$ - -
LGBM $\mathbf{97.78\pm 0.80}$ - -
CB $\mathbf{97.72\pm 0.74}$ - -
These results can be explained as follows. Under the experimental setting adopted from Protocol B, data augmentation is applied prior to dataset splitting, where image flipping generates duplicated samples originating from the same fish instance and retaining identical class labels. Due to the inherent invariance properties of handcrafted descriptors, such transformations produce identical or near-identical feature vectors. As a result, the augmented dataset contains redundant samples with identical handcrafted representations appearing across training and test folds during cross-validation. This redundancy is effectively exploited by models with high learning capacity, such as ANN and tree-based ensembles (RF, ET, LGBM, and CB), leading to exceptionally high performance.
In contrast, linear and distance-based classifiers (LR, SVM, and KNN) do not benefit from sample duplication, as their decision boundaries or neighborhood structures remain mathematically invariant when identical samples are repeated. Consequently, the limited or slightly degraded performance of models indicates that the performance gains observed under Protocol B arise from algorithmic exploitation of data redundancy rather than an increase in the intrinsic discriminative power of the handcrafted features.
5.4.2 Comparative analysis on the non-augmented dataset (Protocol A)
In addition to the replication experiments conducted under Protocol B, we further evaluated our approach under Protocol A using the standard train-test split proposed by Prasetyo et al. [9], who reported a baseline accuracy of 63.21% with their custom deep learning architecture, MB-BE (a modified MobileNetV1), on the original, non-augmented dataset.
@@ 6 Conclusion
This study proposed and systematically evaluated a handcrafted feature engineering approach for fish freshness classification from eye images. We extracted a diverse set of descriptors, including color statistics, variance ratios, percentiles, histograms, and texture features (GLCM and LBP), and integrated them through an incremental fusion strategy to form a rich numerical representation of freshness cues. The extracted features were subsequently used to train and evaluate eight ML models–LR, KNN, SVM, ANN, RF, ET, LGBM, and CB–to comprehensively assess classification performance. Experiments on the FFE dataset demonstrated that, under a standard non-augmented train–test setting, the LGBM model achieved an accuracy of 77.56%, representing an improvement of over 14% compared to the 63.21% reported by Prasetyo et al. [9]. Under an augmented evaluation protocol consistent with prior work, the ANN model obtained 97.49% accuracy, compared to the 77.3% reported by Yildiz et al. [10]. This augmented performance should be interpreted as an upper-bound estimate, while the non-augmented evaluation provides a more realistic proxy for practical deployment. Importantly, comparative analyses between full images and ROI-based segmentation revealed that contextual information from surrounding tissues contributes significantly to classification performance, indicating that discriminative freshness cues extend beyond the eye region alone. Despite the encouraging results obtained under both evaluation protocols, the proposed approach remains laboratory-validated and relies on handcrafted feature engineering, which requires domain expertise and may exhibit limited generalization across unseen species or acquisition conditions. In addition, although handcrafted descriptors demonstrate strong discriminative capability on the FFE dataset, they may fail to capture higher-level abstract patterns that deep learning models can learn automatically when larger and more diverse datasets are available. Future work will therefore focus on hybrid frameworks that combine handcrafted descriptors with deep feature representations, aiming to preserve interpretability while improving robustness and generalization.
@@
Ethical statement:
The authors declare that no ethical approval was required for this study as it exclusively utilizes a publicly available dataset (FFE).
@@
Additional statement:
A preprint version of this manuscript will be uploaded to arXiv for early access and citation.